OpenStudy (anonymous):

A quadratic equation is shown below: 9x2 - 36x + 36 = 0 Part A: Describe the solution(s) to the equation by just determining the discriminant. Show your work. (3 points) Part B: Solve 2x2 - 9x + 7 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used.(4 points)

OpenStudy (anonymous):

9x2 - 36x + 36 = 0 can be written as x2 - 4x + 4 = 0 discriminant = b^2-4*ac= (-4)*(-4) - 4*1*4 =0 so the roots are real and equal

OpenStudy (anonymous):

So the roots are also the solutions?

OpenStudy (anonymous):

for second part factor as 2x^2 - 2x - 7x + 7 = 2x(x-1) -7*(x-1) = (2x-7)(x-1) are the roots or you can use discriminant method also lik a=2, b=-9, c= 7, D = b^2 -4ac = (-9)*(-9) - 4 *2*7 =81-56 =25 so roots are real and unequal

OpenStudy (anonymous):

roots are real and unequal